# On Inverse Problems, Parameter Estimation, and Domain Generalization

### v1

@@ Abstract
Signal restoration and inverse problems are key elements in most real-world data science applications. In the past decades, with the emergence of machine learning methods, inversion of measurements has become a popular step in almost all physical applications, which is normally executed prior to downstream tasks that often involve parameter estimation. In this work, we theoretically analyze the general problem of parameter estimation in an inverse problem setting. First, we address the domain-shift problem by re-formulating it in direct relation with the discrete parameter estimation analysis. We analyze a significant vulnerability in current attempts to enforce domain generalization, which we dubbed the Double Meaning Theorem. Our theoretical findings are experimentally illustrated for domain shift examples in image deblurring and speckle suppression in medical imaging. We then proceed to a theoretical analysis of parameter estimation given observed measurements before and after data processing involving an inversion of the observations. We compare this setting for invertible and non-invertible (degradation) processes, with parameter estimation that is performed directly from given measurements comparing with parameter estimation after data-processing of the observed measurements. We distinguish between continuous and discrete parameter estimation, corresponding with regression and classification problems, respectively. Our theoretical findings align with the well-known information-theoretic data processing inequality, and to a certain degree question the common misconception that data-processing for inversion, based on modern generative models that may often produce outstanding perceptual quality, will necessarily improve the following parameter estimation objective. It is our hope that this paper will provide practitioners with deeper insights that may be leveraged in the future for the development of more efficient and informed strategic system planning, critical in safety-sensitive applications.
Keywords: Cramér-Rao Bound, Data Processing Inequality, Fisher Information, Inverse Problems, Domain Adaptation.
@@ 1 Introduction
Signal restoration and inverse problems have a significant impact on science and engineering applications across numerous disciplines, such as communication, security, healthcare, education and art. As restoration algorithms have a fundamental role in signal processing and imaging sciences, there is a growing need to accurately assess their reliability. Recent advances in generative artificial intelligence had remarkable impact on the ability of inverse-problem solvers to restore corrupted information and to produce perceptually enhanced results. That said, it has been observed that these models are often inclined to generate hallucinations (Alemohammad et al., 2023; Aithal et al., 2024), which in turn has led to further development of uncertainty quantification methods (Vovk et al., 2005; Romano et al., 2019; Angelopoulos et al., 2022) in an attempt to provide statistical guarantees that could hopefully protect from models’ mistakes or hallucinations. Distinct hallucinations were also observed in large language models, e.g., memory distortion, and cognitive mirage, where a model could generate statements that seem reasonable but are either cognitively irrelevant or factually incorrect (Rawte et al., 2024).
The recent popularity of deep learning methods have inspired new interest in the optimal performance bounds in classification (Jeong et al., 2024) and regression tasks
@@
(Shalev-Shwartz and Ben-David, 2014). Machine learning methods have demonstrated outstanding performance for classification tasks. Nevertheless, it is challenging to determine whether such methods have achieved the Bayes error rate, i.e., the lowest error rate obtained by any classifier (Tumer and Ghosh, 1996; Martin et al., 1994). Furthermore, restoration methods (such as denoising, despeckling and deblurring) are increasingly adopted as preprocessing stages prior to downstream tasks, such as classification and segmentation, under the assumption that a restoration pre-processing stage that produces a visually enhanced image, will necessarily improve the results of the following tasks. Whereas, in accordance with the information theoretic data processing inequality (Cover and Thomas, 2006), we know that every processing stage is bound to lose information. It is only possible to extract information and it is impossible to recreate new information.
@@
Most real-world imaging systems acquire measurements that are degraded by noise, scattering, aberrations, attenuation, and smearing. Conventional cameras’ raw sensor measurements are commonly processed in a sequential pipeline, addressing first low-level vision tasks, such as compression, denoising and deblurring, and then proceeding to high-level tasks, such as: classification and feature extraction (Diamond et al., 2021). In medical imaging, it is commonly assumed that image-enhancement pre-processing stages (e.g., denoising, deblurring and super-resolution) would necessarily boost the performance of post-processing parameter-estimation downstream stages. For example, it is often assumed that speckle suppression of ultrasound and optical coherence tomography (OCT) images should be employed prior to parameter estimation, such as segmentation and elasticity assessment to determine the progress of glaucoma (Pereg, 2023b, 2024; Braeu et al., 2024; Yan et al., 2020).
@@
The term class imbalance refers to clinical prediction models with a binary outcome where the event occurrence is much lower than $50\%$ imbalance correction methods, such as: random undersampling, random oversampling, or synthetic minority oversampling technique (SMOTE) led to models with strong miscalibration without better ability to distinguish between patients with and without the outcome event. Clearly, inaccurate probability estimates yield ill-informed clinical decisions, thus reducing the utility of the prediction models. Since the accuracy measure of a classifier depends on the percentage of the true positives and true negatives, a trained classifier may appear to have a high score when the majority of events belong to one class, but the minority class, which is often the high risk class, is misclassified. Similar results were presented for machine learning based prediction models (Carriero et al., 2024).
@@
In the past, most signal restoration algorithms were deterministic, that is, they were designed to provide one restored signal given a degraded input measurement (or several input measurements). Recent stochastic restoration algorithms (Song et al., 2020; Kadkhodaie and Simoncelli, 2021; Ohayon et al., 2023a, b; Delbracio and Milanfar, 2023) explore the possible solutions by sampling from a distribution that is conditioned on the degraded input. This approach applies the inherent nature of most of these ill-conditioned inverse problems, that is - there is more than one solution that can fit the measurement, and there is an inherent uncertainty in real-data applications regarding which solution should be favored. That said, apart from the ability of stochastic approaches to explore different viable solutions, it is still unclear what is their advantage in terms of perceptual quality and robustness. In the context of this work an interesting question is whether stochastic restoration methods can improve the following parameter estimation.
@@
Recent experimental studies observe that image-to-image restoration models are extremely sensitive to adversarial attacks, which in certain cases is diagnosed as mode-collapse. Even visually undetectable input perturbations lead to severe output artifacts (Choi et al., 2019, 2022; Yan et al., 2022). In this work, we establish an alternative reasoning behind this significant vulnerability. We prove that training over several domains in an attempt to improve domain generalization (Double Meaning Theorem (Theorem 3)) may lead to ambiguity in the required output, which in turn inevitably leads to a degraded output. Our theoretical findings are experimentally illustrated for image deblurring and speckle supression.
@@
The general objective of this research work is to provide a mathematical analysis along with the corresponding intuition for the exploration of the optimal path towards parameter estimation in the context of inverse problems. In Section 3, we prove a fundamental property of domain generalization (Double Meaning Theorem (Theorem 3)). We empirically validate our theoretical findings via experiments in natural image deblurring and medical image denoising. In Section 4, we introduce a general problem formulation for parameter estimation in an inverse problem setting, and define a perfect perception estimator. We derive fundamental theorems for Bayes error bounds in classification problems following preprocessing of acquired measurements (Theorems 5-7), and provide illustrative examples to inspire intuition into the proposed point-of-view. Lastly, in Section 5, we prove that in a continuous setting a perfect perception estimator cannot improve the accuracy of the following parameter estimation, whether it is conditioned on that parameter either explicitly or implicitly, but it may achieve equivalent performance (Theorems 10-12).
@@
We note that substantial portion of this paper may be viewed as a review of existing knowledge, and is included for the sake of completeness. We strongly believe that this work emphasizes important aspects for practitioners in real-world applications.
@@ 2.1 Notations
Let $X\in\mathcal{X}^{n\times 1}$ and $Y\in\mathcal{Y}^{d\times 1}$ be random variables that obey stationary and ergodic probability distributions, and have a stationary coupling. Throughout the majority of this work, we consider a setting where $X$ and $Y$ are continuous random variables. It is also possible to consider a practical setting where $X$ and $Y$ are continuous random variables represented in a finite precision machine where both $X$ and $Y$ are quantized into a finite number of discrete values. In the discrete case the proofs generally follow similar outline where the corresponding integrals should be replaced with finite sums. Note that here $Y$ may be a single measurement or a vector of $m\in\mathbb{N}$ measurements. We denote the joint probability of $X$ and $Y$ as $p_{X,Y}(x,y)$, and their corresponding mutual information is defined as $I(X;Y)=\mathcal{D}(p(x,y)||p(x)p(y))\geq 0$, where we have used the standard notation $\mathcal{D}(p||q)\triangleq\int p(u)\log\frac{p(u)}{q(u)}du$ for the Kullback-Liebler (KL) divergence between the probability distributions $p$ and $q$. Namely,
[…]
Equivalently, we assume a sample space that is a set $\Psi$ of paired objects $\Psi=\{\mathbf{y}_{i},\mathbf{x}_{i}\}_{i=1}^{m}$, where $\mathbf{x}_{i}\in\mathcal{X}^{n\times 1}$ are sampled from $p_{\theta}(\mathbf{x})$ and paired with $\mathbf{y}_{i}\in\mathcal{Y}^{d\times 1}$ by a deterministic or stochastic mapping as ground truth. $\Psi_{x}=\{\mathbf{x}_{i}\}_{i=1}^{m}$, can be thought of as possible outcomes of a “perfect” experiment distributed according to an unknown distribution of a certain set of probability measures that depend on the true parameter $\theta$. Whereas $\Psi_{y}=\{\mathbf{y}_{i}\}_{i=1}^{m}$ are the corresponding measured (degraded) outcomes. For example, one the most familiar concrete examples in statistics (Fisher, 1922; Halmos and Savage, 1949) is an $n$-dimensional vector $\mathbf{x}\in\mathbb{R}^{n\times 1}$, of $n$ element-values representing $n$ independent observations of a normally distributed random variable with unknown parameters. The notations $\mathbf{x}$ and $X$ are used below interchangeably.
@@ 2.2 Ergodic Sources
Let $x^{n}$ denote a sequence $x_{1},x_{2},...,x_{n}$, and $\mathbf{x}\in\mathcal{X}^{n\times 1}$ to denote a vector with $n$ entries. In information theory, a stationary stochastic process $u^{n}$ taking values in some finite alphabet $\mathcal{U}$ is referred to as a source. More often than not, communication theory refers to discrete memoryless sources (DMS) (Kramer et al., 2008; Cover and Thomas, 2006). That said, many signals, such as image patches, are usually modeled as entities belonging to some probability distribution forming statistical dependencies (e.g., a Markov random field (MRF) (Roth and Black, 2009; Weiss and Freeman, 2007)) corresponding to the relations between data points in close spatial or temporal proximity. Although the formal definition of ergodic processes is somewhat complicated, as described by Shannon in 1948, the general idea is simple: “In an ergodic process, every sequence that is produced by the process is the same in statistical properties” (Shannon, 1948). Particular sequences generated by the process will exhibit symbol frequencies approaching a definite statistical limit, as the lengths of the sequences is increased. More formally, we assume an ergodic source with memory that emits $n$ symbols from a discrete and finite alphabet $\mathcal{U}$, with probability $P_{U}(u_{1},u_{2},...,u_{n})$. We recall a theorem (Breiman, 1957), here without proof.
Theorem 1 (Entropy and Ergodic Theory (Breiman, 1957))
Let $u_{1},u_{2},...,u_{n}$ be a stationary ergodic process ranging over a finite alphabet $\mathcal{U}$, then there is a constant $H$, defining the entropy rate of the source,
@@
Intuitively, when we observe a source with memory over several time units, once we know the previous source’s entries, the dependencies reduce the overall conditional uncertainty, thus the uncertainty grows more slowly as $n$ grows,. The entropy rate $H$ representing the average uncertainty per time unit, converges over time. Although the entropy rate is typically used for discrete memoryless sources (DMS), every ergodic source has the AEP (McMillan, 1953). The generalization of the Asymptotic Equipartition Property (AEP) to arbitrary ergodic sources can be found in (McMillan, 1953). Entropy typicality applies also to continuous random variables with a density $p_{U}$ replacing the discrete probability $P_{U}(u^{n})$ with the density value $p_{U}(u^{n})$. The AEP leads to Shannon’s source coding theorem stating that the average number of bits required to specify a symbol in a sequence of length $n$, when we consider only the most probable sequences, is $H$. And it is the foundation for the known rate-distortion theory and channel capacity. In Pereg (2023a), we took a similar stand-point with respect to ergodic sources to apply the AEP to bound the sample complexity for supervised learning in regression models with respect to the empirical error and the generalization error.
@@ 2.3 Fisher information
$s(\theta)=\frac{\partial}{\partial\theta}\log p_{\theta}(\mathbf{x}).$ (2)
@@
Fisher information (FI) is defined as the variance of the score (Fisher, 1922),
$J_{X}(\theta)=E\Bigg{\{}\Bigg{(}\frac{\partial}{\partial\theta}\log p_{\theta}(\mathbf{x})\Bigg{)}^{2}\Bigg{|}\theta\Bigg{\}}.$ (3)
We note that $J_{X}(\theta)\geq 0$. If we consider a sample of $m$ random variables $X_{1},X_{2},...,X_{m}$ drawn i.i.d. from $p_{\theta}(x)$, we have $p_{\theta}(x_{1},x_{1},...,x_{m})=\Pi_{i=1}^{m}p_{\theta}(x_{i})$ and
@@
Notably, the chain rule for the FI (Zamir, 1998) is

### v2

@@ Abstract
Signal restoration and inverse problems are key elements in most real-world data science applications. In the past decades, with the emergence of machine learning methods, inversion of measurements has become a popular step in almost all physical applications, normally executed prior to downstream tasks that often involve parameter estimation. In this work, we propose a general framework for theoretical analysis of parameter estimation in inverse problem settings. We distinguish between continuous and discrete parameter estimation, corresponding with regression and classification problems, respectively. We investigate this setting for invertible and non-invertible degradation processes, with parameter estimation that is executed directly from the observed measurements, comparing with parameter estimation after data-processing performing an inversion of the observations. Our theoretical findings align with the well-known information-theoretic data processing inequality, and to a certain degree question the common misconception that data-processing for inversion, based on modern generative models that may often produce outstanding perceptual quality, will necessarily improve the following parameter estimation objective. Importantly, by re-formulating the domain-shift problem in direct relation with discrete parameter estimation, we expose a significant vulnerability in current popular practical attempts to enforce domain generalization, which we dubbed the Double Meaning Theorem. These theoretical findings are experimentally illustrated for domain shift examples in image deblurring and speckle suppression in medical imaging. It is our hope that this paper will provide practitioners with deeper insights that may be leveraged in the future for the development of more efficient and informed strategic system planning, critical in safety-sensitive applications.
keywords:
Cramér-Rao Bound, Data Processing Inequality, Fisher Information, Inverse Problems, Domain Adaptation.
{MSCcodes}
68Q25, 68R10, 68U05
@@ 1 Introduction
Signal restoration and inverse problems have a significant impact on science and engineering applications across numerous disciplines, such as communication, security, healthcare, education and art. As restoration algorithms have a fundamental role in signal processing and imaging sciences, there is a growing need to accurately assess their reliability. Recent advances in generative artificial intelligence had remarkable impact on the ability of inverse-problem solvers to restore corrupted information and to produce perceptually enhanced results. That said, it has been observed that these models are often inclined to generate hallucinations [Alemohammad:2023, Aithal:2024], which in turn has led to further development of uncertainty quantification methods [Vovk:2005, Romano:2019, Angelopoulos:2022] in an attempt to provide statistical guarantees that could hopefully protect from models’ mistakes or hallucinations. Distinct hallucinations were also observed in large language models, e.g., memory distortion, and cognitive mirage, where a model could generate statements that seem reasonable but are either cognitively irrelevant or factually incorrect [Rawte:2024].
@@
The recent popularity of deep learning methods have inspired new interest in optimal performance bounds in classification [Jeong:2024] and regression tasks [Shalev:2014]. Machine learning methods have demonstrated outstanding performance for classification tasks. Nevertheless, it is challenging to determine whether such methods have achieved the Bayes error rate, i.e., the lowest error rate obtained by any classifier [Tumer:1996, Martin:1994]. Furthermore, restoration methods (such as denoising, despeckling and deblurring) are increasingly adopted as preprocessing stages prior to downstream tasks, such as classification and segmentation, under the assumption that a restoration pre-processing stage that produces a visually enhanced image, will necessarily improve the results of the following tasks. Whereas, in accordance with the information theoretic data processing inequality [ThomasCover:2006], we know that every processing stage is bound to lose information. It is only possible to extract information and it is impossible to recreate new information.
@@
Most real-world imaging systems acquire measurements that are degraded by noise, scattering, aberrations, attenuation, and smearing. Conventional cameras’ raw sensor measurements are commonly processed in a sequential pipeline, addressing first low-level vision tasks, such as compression, denoising and deblurring, and then proceeding to high-level tasks, such as: classification and feature extraction [Diamond:2021]. In medical imaging, it is commonly assumed that image-enhancement pre-processing stages (e.g., denoising, deblurring and super-resolution) would necessarily boost the performance of post-processing parameter-estimation downstream stages. For example, it is often assumed that speckle suppression of ultrasound and optical coherence tomography (OCT) images should be employed prior to parameter estimation, such as segmentation and elasticity assessment to determine the progress of glaucoma [Pereg:2023B, Pereg:2024A, Braeu:2024, Yan:2020].
@@
The term class imbalance refers to clinical prediction models with a binary outcome where the event occurrence is much lower than $50\%$ imbalance correction methods, such as: random undersampling, random oversampling, or synthetic minority oversampling technique (SMOTE) led to models with strong miscalibration without better ability to distinguish between patients with and without the outcome event. Clearly, inaccurate probability estimates yield ill-informed clinical decisions, thus reducing the utility of the prediction models. Since the accuracy measure of a classifier depends on the percentage of the true positives and true negatives, a trained classifier may appear to have a high score when the majority of events belong to one class, but the minority class, which is often the high risk class, is misclassified. Similar results were presented for machine learning based prediction models [Carriero:2024].
@@
In the past, most signal restoration algorithms were deterministic, that is, they were designed to provide one restored signal given a degraded input measurement (or several input measurements). Recent stochastic restoration algorithms [Song:2020, Simoncelli:2021, Ohayon:2023a, Ohayon:2023b, Milanfar:2023] explore the possible solutions by sampling from a distribution that is conditioned on the degraded input. This approach applies the inherent nature of most of these ill-conditioned inverse problems, that is - there is more than one solution that can fit the measurement, and there is an inherent uncertainty in real-data applications regarding which solution should be favored. That said, apart from the ability of stochastic approaches to explore different viable solutions, it is still unclear what is their advantage in terms of perceptual quality and robustness. In the context of this work an interesting question is whether stochastic restoration methods can improve the following parameter estimation.
@@
Recent experimental studies observe that image-to-image restoration models are extremely sensitive to adversarial attacks, which in certain cases is diagnosed as mode-collapse. Namely, even visually undetectable input perturbations lead to severe output artifacts [Choi:2019, Choi:2022, Yan:2022]. In this work, we establish an alternative reasoning behind this significant vulnerability. We prove that training over several domains in an attempt to improve domain generalization (Double Meaning Theorem (Theorem 3.2)) may lead to ambiguity in the required output, which in turn inevitably leads to a degraded output. Our theoretical findings are experimentally illustrated for image deblurring and speckle supression.
@@
The general objective of this research work is to provide a mathematical analysis along with the corresponding intuition for the exploration of the optimal path towards parameter estimation in the context of inverse problems. To this end, we analyze several specific settings. In Section 3, we prove a fundamental property of domain generalization (Double Meaning Theorem (Theorem 3.2)). We empirically validate our theoretical findings via experiments in natural image deblurring and medical image denoising. In Section 4, we introduce a general problem formulation for parameter estimation in an inverse problem setting, and define a perfect perception estimator. We derive fundamental theorems for Bayes error bounds in classification problems following preprocessing of acquired measurements (Theorems 4.2-4.4), and provide illustrative examples to inspire intuition into the proposed point-of-view. Lastly, in Section 5, we prove that in a continuous setting a perfect perception estimator cannot improve the accuracy of the following parameter estimation, whether it is conditioned on that parameter either explicitly or implicitly, but it may achieve equivalent performance (Theorems 5.3-5.7).
@@
We note that a substantial portion of this paper may be viewed as a review of existing knowledge, and is included for the sake of completeness. We strongly believe that this work emphasizes important aspects for practitioners in real-world applications.
@@ 2.1 Notations
Let $X\in\mathcal{X}^{n\times 1}$ and $Y\in\mathcal{Y}^{d\times 1}$ be random variables that obey stationary and ergodic probability distributions, and have a stationary coupling [Gray:2009]. Throughout the majority of this work, we consider a setting where $X$ and $Y$ are continuous random variables. It is also possible to consider a practical setting where $X$ and $Y$ are continuous random variables represented in a finite precision machine where both $X$ and $Y$ are quantized into a finite number of discrete values. In the discrete case the proofs generally follow similar outline where the corresponding integrals should be replaced with finite sums. Note that here $Y$ may be a single measurement or a vector of $m\in\mathbb{N}$ measurements. We denote the joint probability of $X$ and $Y$ as $p_{X,Y}(x,y)$, and their corresponding mutual information is defined as $I(X;Y)=\mathcal{D}(p(x,y)||p(x)p(y))\geq 0$, where we have used the standard notation $\mathcal{D}(p||q)\triangleq\int p(u)\log\frac{p(u)}{q(u)}du$ for the Kullback-Liebler (KL) divergence between the probability density functions $p$ and $q$. Namely, in a continuous setting,
(1) $I(X;Y)=\int p(x,y)\mathrm{log}\frac{p(x,y)}{p(x)p(y)}dx\ dy.$
Equivalently, we assume a sample space that is a set $\Psi$ of paired objects $\Psi=\{\mathbf{y}_{i},\mathbf{x}_{i}\}_{i=1}^{m}$, where $\mathbf{x}_{i}\in\mathcal{X}^{n\times 1}$ are sampled from $p_{\theta}(\mathbf{x})$ and paired with $\mathbf{y}_{i}\in\mathcal{Y}^{d\times 1}$ by a deterministic or stochastic mapping as ground truth. $\Psi_{x}=\{\mathbf{x}_{i}\}_{i=1}^{m}$, can be thought of as possible outcomes of a “perfect” experiment distributed according to an unknown distribution of a certain set of probability measures that depend on the true parameter $\theta$. Whereas $\Psi_{y}=\{\mathbf{y}_{i}\}_{i=1}^{m}$ are the corresponding measured (degraded) outcomes. For example, one the most familiar concrete examples in statistics [Fisher:1922, Halmos:1949] is an $n$-dimensional vector $\mathbf{x}\in\mathbb{R}^{n\times 1}$, of $n$ element-values representing $n$ independent observations of a normally distributed random variable with unknown parameters. The notations $\mathbf{x}$ and $X$ are used below interchangeably.
@@ 2.2 Ergodic Sources
Let $x^{n}$ denote a sequence $x_{1},x_{2},...,x_{n}$, and $\mathbf{x}\in\mathcal{X}^{n\times 1}$ to denote a vector with $n$ entries. In information theory, a stationary stochastic process $u^{n}$ taking values in some finite alphabet $\mathcal{U}$ is referred to as a source. More often than not, communication theory refers to discrete memoryless sources (DMS) [Kramer:2008, ThomasCover:2006]. That said, many signals, such as image patches, are usually modeled as entities belonging to some probability distribution forming statistical dependencies (e.g., a Markov random field (MRF) [Roth:2009, Weiss:2007]) corresponding to the relations between data points in close spatial or temporal proximity. Although the formal definition of ergodic processes is somewhat complicated ([Gray:2009, Papoulis:1991], Appendix A), as described by Shannon in 1948, the general idea is simple: “In an ergodic process, every sequence that is produced by the process is the same in statistical properties” [Shannon:1948]. Particular sequences generated by the process will exhibit symbol frequencies approaching a definite statistical limit, as the lengths of the sequences is increased. More formally, we assume an ergodic source with memory that emits $n$ symbols from a discrete and finite alphabet $\mathcal{U}$, with probability $P_{U}(u_{1},u_{2},...,u_{n})$. We recall a theorem [Breiman:1957], here without proof.
Theorem 2.1 (Entropy and Ergodic Theory [Breiman:1957]).
Let $(u_{n})_{n\in\mathbb{Z}}$ be a stationary ergodic process ranging over a finite alphabet $\mathcal{U}$, then there is a constant $H$, defining the entropy rate of the source,
@@
Intuitively, when we observe a source with memory over several time units, once we know the previous source’s entries, the dependencies reduce the overall conditional uncertainty, thus the uncertainty grows more slowly as $n$ grows. The entropy rate $H$ representing the average uncertainty per time unit, converges over time. Although the entropy rate is typically used for discrete memoryless sources (DMS), every ergodic source has the Asymptotic Equipartition Property (AEP) [Breiman:1953]. The generalization of the AEP to arbitrary ergodic sources can be found in [Breiman:1953]. Entropy typicality applies also to continuous random variables with a density $p_{U}$ replacing the discrete probability $P_{U}(u^{n})$ with the density value $p_{U}(u^{n})$. The AEP leads to Shannon’s source coding theorem stating that the average number of bits required to specify a symbol in a sequence of length $n$, when we consider only the most probable sequences, is $H$. And it is the foundation for the known rate-distortion theory and channel capacity. In Pereg et al. (2023) [Pereg:2023A], we took a similar stand-point with respect to ergodic sources to apply the AEP to bound the sample complexity for supervised learning in regression models with respect to the empirical error and the generalization error.
@@ 2.3 Fisher information
(2) $s(\theta)=\frac{\partial}{\partial\theta}\log p_{\theta}(\mathbf{x}).$
@@
Fisher information (FI) is defined as the variance of the score [Fisher:1922],
(3) $J_{X}(\theta)=E\Bigg\{\Bigg(\frac{\partial}{\partial\theta}\log p_{\theta}(\mathbf{x})\Bigg)^{2}\Bigg|\theta\Bigg\}.$
$J_{X}(\theta)\geq 0$. The FI is defined only under regulatory conditions of differentiability and integrability. For example, a simple case where the Cramér-Rao bound may not be valid is when the support of $p_{\theta}(\mathbf{x})$ depends on $\theta$. If we consider a sample of $m$ random variables $X_{1},X_{2},...,X_{m}$ drawn i.i.d. from $p_{\theta}(x)$, we have $p_{\theta}(x_{1},x_{1},...,x_{m})=\Pi_{i=1}^{m}p_{\theta}(x_{i})$ and
@@
Notably, the chain rule for the FI [Zamir:1998] is